Suppose there is a market demand fuction given by:p=30-1/2Qd,and market supply function :Qs=30+2p,what would be equilibrium price and quantity (respectively)



1
Expert's answer
2022-01-25T06:47:08-0500

"P=30-\frac{1}{2}Q_d"

"\frac{1}{2}Q_d=30-P"

"Q_d= 60-2P"

"Q_s=30+2P"

At Equilibrium, "Q_s=Q_d"

"60-2P=30+2P"

"4P=30"

P= 7.5

Replace P to either demand or Supply equation

"Q_s= 30+2(7.5)= 30+15"

Equilibrium Quantity= 45